OpenStudy (anonymous):

please help!!!!!!!!! Find the area of a square tile if the width is \sqrt {0.5} in.

OpenStudy (blurbendy):

w = \[\sqrt{0.5}\] ?

OpenStudy (anonymous):

yeah You have to find the area

OpenStudy (blurbendy):

square means all side lengths are the same A = sidelength^2 A = \[(\sqrt{0.5})^2\]

OpenStudy (anonymous):

Thats the answer ?

OpenStudy (blurbendy):

simplify it

OpenStudy (anonymous):

i got .25

OpenStudy (blurbendy):

you should get 1/2

OpenStudy (anonymous):

I just plugged it in the calc

OpenStudy (blurbendy):

if you ever square something inside a square root, you will always get what's inside the square root back

OpenStudy (anonymous):

Find the area of a square patio if the length is \sqrt {250} ft. so what would this one be ?

OpenStudy (blurbendy):

same idea as the last question

OpenStudy (anonymous):

ok but if i sqaure 0.5 I get .25

OpenStudy (anonymous):

ohhhhhh i see but how'd you know it was 1/2 ?

OpenStudy (blurbendy):

because it's a math rule if you square something inside of a square root you ALWAYS get what's inside the square root back

OpenStudy (anonymous):

so the other one would be 250

OpenStudy (blurbendy):

yes

OpenStudy (anonymous):

Find the area of a rectangular basketball court if the length is \sqrt {8839} ft. and the width is \sqrt {2500} ft.

OpenStudy (anonymous):

what about this ?

OpenStudy (blurbendy):

area of rectangle = l * w

OpenStudy (anonymous):

so you multiply them